D2y/dx2-24dy/dx+144=0


1
Expert's answer
2021-05-20T07:40:02-0400

"\frac{D^2y}{dx^2}-24\frac{dy}{dx}+144=0"


the differential equation is of the form

"y''+py'+qy=s"


p=-24

q=0

s=-144

It is linear inhomogeneous

second -order differential equation with constant coefficients.

the equation has an easy solution

first we should find the roots of the characteristic equation

"q+(k^2+kp)=0"

"k^2+kp=0"

this is simple equation

the roots of this equation are


"k_1=0 \\
k_2=24"

as there are two roots of characteristic equation ,

and the roots are not complex , then

solving the corresponding differential equation looks as follows


"\implies"

"y(x)=C_1e^{k_1x}+C_2e^{k_2x}"

"y(x)=C_1+C_2e^{24x}"


we get asolution for the corresponding homogeneous equation

Now we should solve the inhomogeneous equation


y''+py'+qy=s


use variation of parameters method

Suppose that C1 and C2 - it is function of x


tjhe general solution is

"y(x)=C_1(x)+C_2(x)e^{24x}"


where C1(x) and C2(x)

by the method of variation of parameters , we find the solution from the system:

"y_1(x)\frac{d}{dx}C_1(x)+y_2(x)\frac{d}{dx_2}C_2(x)=0\\"


"\frac{d}{dx}C_1(x)\frac{d}{dx}y_1(x)+\frac{d}{dx}C_2(x)\frac{d}{dx}y_2(x)=f(x)"


where

y1(x) and y2(x) - linearly independent particular solutions of Linear Ordinary Differential Equations,

y1(x)=1 (C1=1 , C2=0),

y2(x)="e^{(24x)}" (C1=0 , C2=1)

the free term f=s or

"f(x)=-144"

so the system has the form


"\\e^{24x}\frac{d}{dx}C_2(x)+\frac{d}{dx}C_1(x)=0\\
\frac{d}{dx}1\frac{d}{dx}C_1(x)+\frac{d}{dx}C_2(x)\frac{d}{dx}e^{24x}=-144 \\
or \\
e^{24x}\frac{d}{dx}C_2(x)+\frac{d}{dx}C_1(x)=0 \\
24e^{24x}\frac{d}{dx}C_2(x)=-144"


solve the system :

"\frac{d}{dx}C_1(x)=6
\\
\frac{d}{dx}C_2(x)=-6e^{-24x}"

it is the simple equation , solve these equation

"C_1(x)=C_3+ \int6dx\\
C_2(x)=C_4+\int(-6e^{-24x})dx"


or

"C_1(x)=C_3+6x \\
C_2(x)=C_4+\frac{e^{-24}}{4}"


substitute found C1(x) and C2(x) to

"y(x)=C_1(x)+C_2(x)e^{24x}"


the answer is

"y(x)=C_3+C_4e^{24x}+6x+\frac{1}{4}"


where C3 and C4 are constant and also"\frac{1}{4}"


so final answer is


"\boxed{y{\left(x \right)} = C_{1} + C_{2} e^{24 x} + 6 x}answer"